A 0.20 Ω bulb is connected to a 12-volt power supply. The power used by an element in a circuit can be calculated with P = I2 R , where I is the current through the element and R is the resistance of the element.

(a) What is the power delivered to the bulb in W?

(b) What is the total power delivered by the battery in W?
Answer #1

I = V/R = 12 V / 0.20 ohm = 60 A

(A) P delivered = I^2 R = (60^2)(0.20) = 720 W

(B) by battery = V I = (12 V) (60 A) = 720 W